What is the solution set to the following equations: 3x-2y=4,4x+5y=28
Answer by stanbon(75887)

What is the solution set to the following equations:
3x-2y=4
4x+5y=28
----------
Multiply thru the 1st equation by 5; Multiply thru the 2nd by 2
---------
15x-10y=20
8x+10y=56
---------------
Add the two equations and solve for "x":
23x = 76
x = 76/23
---------------------
Substitute into 3x-2y = 4 to solve for "y":
3(76/23) - 2y = 4
(228/23) - 2y = (92/23)
2y = 136/23
y = 68/23
====================
Cheers,
Stan H.